# Enhancement of axial anomaly effects in hot two-color QCD: FRG approach in the linear sigma model

### v1

@@ Abstract
We investigate the thermal properties of hadrons in two-color quantum chromodynamics (QC2D) using the functional renormalization group (FRG) method, with particular focus on modifications of the $U(1)$ axial anomaly effects. The hadrons are described by a linear sigma model (LSM) based on the Pauli-Gürsey $SU(4)$ symmetry, which incorporates both low-lying $0^{\pm}$ mesons and diquark baryons. We find that all quartic couplings are comparably suppressed when physical values of the pion mass and decay constant are taken as inputs, for which a reasonably smooth chiral symmetry restoration at finite temperature is reproduced. Consequently, mass differences among chiral partners remain small. Despite these tiny mass differences, mass degeneracies of chiral partners in the hot medium are clearly demonstrated, consistent with chiral symmetry restoration. Moreover, we find that the couplings responsible for the $U(1)$ axial anomaly are enhanced upon entering the finite temperature regime. Baryonic fluctuations also provide sizable contributions to these enhancements. Finally, the fate of the topological susceptibility in the hot QC2D medium is examined.
@@ I Introduction
The $U(1)$ axial anomaly, being responsible for the origin of a large $\eta^{\prime}$ meson mass, is one of the most important symmetry properties of quantum chromodynamics (QCD). Unlike the breakdown of chiral symmetry, the $U(1)_{A}$ anomaly does not arise from spontaneous symmetry breaking, but caused by topological fluctuations of gluons, i.e., instantons ’t Hooft (1976a). Delineating the effects of the $U(1)_{A}$ anomaly in the hadronic regime helps illuminate interactions that cannot be fully explained by chiral symmetry alone.
@@
In the hadronic regime, within chiral effective models, such as the linear sigma model (LSM), the $U(1)_{A}$ anomaly effects can be well incorporated by the so-called Kobayashi-Maskawa-’t Hooft (KMT) determinant Kobayashi and Maskawa (1970); Kobayashi et al. (1971); ’t Hooft (1976b, a). The coefficient of the KMT term is proportional to the instanton density in the underlying theory, according to a weak coupling calculation on an instanton background. Thus, one can easily predict the disappearance of the anomaly, and therefore the determinant contribution of KMT in the effective model, at the high temperature and/or density limit, due to the Debye screening of electric gluons Gross et al. (1981); Rapp et al. (2000). At intermediate temperature ranges, on the other hand, the $U(1)_{A}$ anomaly effects on hadrons are not well understood Fejos (2015).
@@
In Refs. Fejos and Hosaka (2016); Fejos and Patkos (2022), based on a functional renormalization group (FRG) treatment for the $N_{f}=2+1$ LSM, a scenario where the $U(1)_{A}$ anomaly effects are strengthened at finite temperature due to mesonic fluctuations was proposed. Motivated by this finding, in this paper, we investigate the fate of the anomaly effects at finite temperature in two-color QCD (QC2D), in order to theoretically confirm whether the anomaly enhancement is universal to the number of colors. In what follows, we apply the FRG method to a LSM of QC2D with two flavors ($N_{f}=2$), which was introduced in Ref. Suenaga et al. (2023). This version of the LSM has been developed mainly to investigate the cold and dense properties of QC2D matter extensively Kawaguchi and Suenaga (2023, 2024); Suenaga et al. (2024) (for a review, see Suenaga (2025)).
@@
The FRG, as a mathematical tool, is one of the most promising functional methods in quantum field theory, capable of capturing essential features of strongly interacting quantum systems Berges et al. (2002); Dupuis et al. (2021). To the best of our knowledge, there have been no attempts using the FRG method to describe the thermodynamics of QC2D with both mesonic and baryonic degrees of freedom included. In addition to exploring the thermal behavior of the $U(1)_{A}$ anomaly, this paper also aims to establish an FRG-based framework capable of accounting for the thermodynamic properties of two-color QCD in low energy regimes Fejos and Suenaga (2025).
@@
In QC2D, (anti)diquarks composed of two (anti)quarks are counted as color singlet hadrons. These (anti)diquark baryons can be treated on an equal footing with mesons due to the pseudoreality of the color group $SU(2)$: ${\bm{2}}\simeq\bar{\bm{2}}$. In terms of chiral representation, this property is reflected by an extension of chiral symmetry; in QC2D, $SU(N_{f})_{L}\times SU(N_{f})_{R}$ chiral symmetry is extended to an $SU(2N_{f})$ one, where (anti)diquark baryons and mesons belong to the same multiplet Kogut et al. (1999, 2000). This extended symmetry is often referred to as the Pauli-Gürsey $SU(2N_{f})$ symmetry Pauli (1957); Gürsey (1958).
@@
Since the LSM in QC2D is constructed upon the Pauli-Gürsey symmetry, one can naturally incorporate baryonic contributions in addition to mesonic ones in evaluating the $U(1)_{A}$ anomaly effects. The diquark contributions cannot be easily captured in ordinary three-color QCD, and thus we will pay particular attention to separating baryonic and mesonic contributions in order to gain insights into fluctuations specific to QC2D.
@@
In-medium properties of QC2D are currently being investigated by first-principles lattice simulations, especially due to the absence of the sign problem at finite baryon chemical potential Boz et al. (2020); Buividovich et al. (2020); Astrakhantsev et al. (2020); Iida et al. (2024); Braguta (2023). Therefore, in the future, one may expect numerical experiments on the lattice, which are capable of testing our current findings. In QC2D, there is no a priori physical point. In the present analysis, we take physical values of the pion mass and the pion decay constant, while $\eta$ mass is considered as a free parameter, in order to investigate the $U(1)_{A}$ anomaly effects in a wider range. We also examine the fate of the anomaly effects, as well as hadron mass modifications in a hot medium to provide useful information for future numerical experiments.
@@
This article is organized as follows. In Sec. II we briefly review the emergence of the Pauli-Gürsey $SU(4)$ symmetry in QC2D with $N_{f}=2$ and introduce our LSM. In Sec. III our approach, based on the FRG framework, is explained, and RG flow equations for the couplings are derived. Then, in Sec. IV we present numerical results on the anomaly coefficients, hadron masses, and the topological susceptibility at finite temperature, by solving the flow equations. Section V is devoted to detailed analyses of baryonic contributions. Finally, Sec. VI is devoted to the conclusions.
@@ II Two-flavor LSM in QC2D
In this section we briefly review the two-flavor LSM in QC2D introduced in Ref. Suenaga et al. (2023).
@@
${\cal L}_{\rm QC_{2}D}$ $=$ $\bar{\psi}i{\ooalign{\hfil/\hfil\crcr$D$}}\psi$ (3)
@@
where in the second line we made use of the pseudorealities $\tau_{c}^{a}=\tau_{c}^{2}(\tau_{c}^{a})^{T}\tau_{c}^{2}$ and $\sigma^{i}=-\sigma^{2}(\sigma^{i})^{T}\sigma^{2}$ Kogut et al. (1999, 2000). In Eq. (4), the covariant derivative reads $D_{\mu}={\partial}_{\mu}-ig_{s}{A}_{\mu}^{a}\tau_{c}^{a}$ with the strong coupling constant $g_{s}$ and $\sigma^{\mu}=({\bm{1}},\sigma^{i})$. Intuitively speaking, the second line of Eq. (4) implies that gluons cannot distinguish quarks and antiquarks due to the pseudoreality of the $SU(2)_{c}$ group and they can be described by a single multiplet $\Psi$ collectively. As a result, the Lagrangian with massless quarks possesses the so-called Pauli-Gürsey $SU(4)$ symmetry, as an extension of $SU(2)_{L}\times SU(2)_{R}$ chiral symmetry Pauli (1957); Gürsey (1958).
Chiral effective models in QC2D must respect the Pauli-Gürsey $SU(4)$ symmetry in order to accurately describe low-energy hadronic excitations. In particular, the LSM is based on the linear representation of this $SU(4)$ symmetry, in which hadron fields are arranged into a $4\times 4$ matrix $\Sigma$. The transformation of $\Sigma$ under the $SU(4)$ group is given by an underlying quark bilinear: $\Sigma_{ij}\sim\Psi^{T}_{j}\sigma^{2}\tau_{c}^{2}\Psi_{i}$ ($i,j=1$ $4$). Hence, $\Sigma$ transforms as
@@
where ${\cal S}^{a}$ and ${\cal P}^{a}$ ($a=0$ $5$) are defined through the meson fields
@@
Here $B$ ($\bar{B}$) and $B^{\prime}$ ($\bar{B}^{\prime}$) are $0^{+}$ and $0^{-}$ (anti)diquark baryons, respectively. We note that mesons ($\bar{q}q$) and baryons ($qq$) are treated in a unified way reflecting the pseudoreality of $SU(2)$ color group: ${\bm{2}}\simeq\bar{\bm{2}}$. In Eq. (6), $4\times 4$ matrices $X^{a}$ are generators belonging to the algebra of $G/H=SU(4)/Sp(4)$,
$X^{a=1-3}$ $=$ $\frac{1}{2\sqrt{2}}\left(\begin{array}[]{cc}\tau_{f}^{a}&0\\ 0&(\tau_{f}^{a})^{T}\\ \end{array}\right)\ ,$ (11)
@@
where $\tau_{f}^{a}$ ($a=1$ $3$) are the Pauli matrices acting on flavor space, and $D^{a=4}=\tau_{f}^{2}$ and $D_{f}^{a=5}=i\tau_{f}^{2}$. Furthermore, $X^{0}={\bm{1}}_{4\times 4}/(2\sqrt{2})$, and
@@
It should be noted that in QC2D with $N_{f}=2$, the breaking pattern of chiral symmetry is $SU(4)\to Sp(4)$, leading to appearance of five Nambu-Goldstone (NG) bosons: three pions, a diquark baryon, and an antidiquark baryon.
With Eq. (6), the Lagrangian of LSM takes the form of Suenaga et al. (2023); Suenaga (2025)
@@
where $I_{A}$ is responsible for the $U(1)_{A}$ anomaly, since $\Sigma\overset{U(1)_{A}}{\to}{\rm e}^{-i\theta_{A}}\Sigma{\rm e}^{-i\theta_{A}}={\rm e}^{-2i\theta_{A}}\Sigma$ [^1] The terms containing $\lambda_{1}$ and $\lambda_{2}$ are rearranged as $\lambda_{1}I_{1}^{2}+\lambda_{2}I_{2}=\left(\lambda_{1}-\frac{\lambda_{2}}{4}\right)\big{(}{\rm tr}[\Sigma^{\dagger}\Sigma]\big{)}^{2}+\lambda_{2}{\rm tr}[(\Sigma^{\dagger}\Sigma)^{2}]\ .$ Thus, matching with the standard notation $\bar{\lambda}_{1}\big{(}{\rm tr}[\Sigma^{\dagger}\Sigma]\big{)}^{2}+\bar{\lambda}_{2}{\rm tr}[(\Sigma^{\dagger}\Sigma)^{2}]$ yields $\bar{\lambda}_{1}=\lambda_{1}-\lambda_{2}/4$ and $\bar{\lambda}_{2}=\lambda_{2}$. [^2] We note that the determinant term is not an independent invariant, due to the $\det\Sigma+\det\Sigma^{\dagger}=\frac{1}{4}I_{A}^{2}+\frac{1}{2}I_{2}-\frac{1}{8}I_{1}^{2}$ relation, therefore, it must be omitted from the potential..
@@ III.1 Basics
In this section, we introduce the concept of the FRG method and derive flow equations for the couplings of the LSM to incorporate quantum and thermal fluctuations. We do not intend to give a comprehensive description on the technique; the reader is referred to the following reviews Berges et al. (2002); Dupuis et al. (2021).
The core of the FRG formalism lies on the scale dependent quantum effective action, $\Gamma_{k}$, which, unlike the ordinary quantum effective action, $\Gamma$, incorporates fluctuations only with momenta $q\gtrsim k$, where $k$ plays the role of a scale separation variable. This construction is achieved by introducing a momentum dependent mass term into the classical action, usually called the regulator, which gives a large mass to the IR modes, effectively freezing and thus excluding them from the path integral. As $k\rightarrow 0$, one demands that the regulator vanishes, leading $\Gamma_{k=0}=\Gamma$, while for $k\rightarrow\infty$ (or practically some UV cutoff $\Lambda_{\rm UV}$) requires that the regulator tends to infinity for all modes, realizing that $\Gamma_{k\rightarrow\infty}=S$, $S$ being the classical action, upon which our quantum field theory is built.
The scale dependent effective action, $\Gamma_{k}$, obeys the Wetterich equation Wetterich (1993):
@@
~$\tilde{\partial}_{k}=\partial_{k}R_{k}\frac{\partial_{k}}{\partial R_{k}},$ (24)
@@
where once again, both the Tr and Log operations need to be taken in the functional and matrix senses. One may also say that ~$\tilde{\partial}_{k}$ is a differential operator that only acts on the $k$ dependence of $R_{k}$.
The right-hand side (RHS) of Eq. (23) or Eq. (25) is generally nonlocal since $\Gamma_{k}^{\prime\prime}$ remains as a function of ${\cal S}^{a}(x)$ and ${\cal P}^{a}(x)$. In the present analysis, however, we employ the local potential approximation (LPA) and ignore nonlocalities. In the LPA the scale dependent effective action is approximated as ($\int_{x}\equiv\int d^{4}x_{E}$)
$\Gamma_{k}=\int_{x}\Big{[}{\rm tr}[\partial_{i}\Sigma^{\dagger}\partial_{i}\Sigma]+c{\rm tr}[\zeta^{\dagger}\Sigma+\Sigma^{\dagger}\zeta]+V_{k}\Big{]},$ (26)
where $V_{k}$ is now a local function, called the scale dependent effective potential. Note that, we have adopted the Euclidean signature compared to Eq. (18). Also note that, the explicit symmetry breaking term does not depend on the separation scale $k$, which follows from the fact that couplings corresponding to linear terms in the field variable do not renormalize. By employing the LPA ansatz (26), what we need to establish is a flow equation for $V_{k}$.
@@
$\Gamma_{k}^{\prime\prime}(p,q)=\big{(}p^{2}+V_{k}^{\prime\prime}+R_{k}(p)\big{)}\delta(p+q),$ (27)
where the momentum variables contain both the frequencies and the three-momenta, therefore, in accordance, $p^{2}=\omega_{n}^{2}+{\bm{p}}^{2}$ with $\omega_{n}=2\pi nT$ ($n\in{\mathbb{Z}}$) being the Matsubara frequencies. In the present analysis we adopt the so-called 3D Litim regulator Litim (2001), which, in Fourier space, depends only on the three-momentum:
@@
where now both the trace and log operations need to be considered only in the matrix sense, and the RHS includes one single momentum integral only. Note that, owing to the LPA, wave function renormalizations of all field variables are omitted. The summation over $n$ stands for the infinite sum of the Matsubara frequencies $\omega_{n}$, explicitly showing that we are working at some finite temperature $T$. At the UV scale $k=\Lambda_{\rm UV}$, and $V_{k=\Lambda_{\rm UV}}$ is reduced to the classical potential (22), while $V_{k=0}$ becomes the full quantum effective potential. In solving the flow equation, we need to integrate from $k=\Lambda_{\rm UV}$ to $k=0$ using the initial condition dictated by Eq. (22).

### v2

@@ Abstract
We investigate the thermal properties of hadrons in two-color quantum chromodynamics (QC2D) using the functional renormalization group method, with particular focus on modifications of the $U(1)$ axial anomaly effects. The hadrons are described by a linear sigma model based on the Pauli-Gürsey $SU(4)$ symmetry, which incorporates both low-lying $0^{\pm}$ mesons and diquark baryons. We find that all quartic couplings are comparably suppressed when physical values of the pion mass and decay constant are taken as inputs, for which a reasonably smooth chiral symmetry restoration at finite temperature is reproduced. Consequently, mass differences among chiral partners remain small. Despite these tiny mass differences, mass degeneracies of chiral partners in the hot medium are clearly demonstrated, consistent with chiral symmetry restoration. Moreover, we find that the couplings responsible for the $U(1)$ axial anomaly are enhanced upon entering the finite-temperature regime. Baryonic fluctuations also provide sizable contributions to these enhancements. Finally, the fate of the topological susceptibility in the hot QC2D medium is examined.
@@ I Introduction
The $U(1)$ axial anomaly, being responsible for the origin of a large $\eta^{\prime}$ meson mass, is one of the most important symmetry properties of quantum chromodynamics (QCD). Unlike the breakdown of chiral symmetry, the $U(1)_{A}$ anomaly does not arise from spontaneous symmetry breaking but is caused by topological fluctuations of gluons, i.e., instantons tHooft:1976rip. Delineating the effects of the $U(1)_{A}$ anomaly in the hadronic regime helps illuminate interactions that cannot be fully explained by chiral symmetry alone.
@@
In the hadronic regime, within chiral effective models, such as the linear sigma model (LSM), the $U(1)_{A}$ anomaly effects can be well incorporated by the so-called Kobayashi-Maskawa–’t Hooft (KMT) determinant Kobayashi:1970ji; Kobayashi:1971qz; tHooft:1976snw; tHooft:1976rip. The coefficient of the KMT term is proportional to the instanton density in the underlying theory, according to a weak coupling calculation on an instanton background. Thus, one can easily predict the disappearance of the anomaly, and, therefore the determinant contribution of KMT in the effective model, at the high temperature and/or density limit, due to the Debye screening of electric gluons Gross:1980br; Rapp:1999qa. At intermediate temperature ranges, on the other hand, the $U(1)_{A}$ anomaly effects on hadrons are not well understood Fejos:2015xca.
@@
In Refs. Fejos:2016hbp; Fejos:2021yod, based on a functional renormalization group (FRG) treatment for the $N_{f}=2+1$ LSM, a scenario where the $U(1)_{A}$ anomaly effects are strengthened at finite temperature due to mesonic fluctuations was proposed. Motivated by this finding, in this paper, we investigate the fate of the anomaly effects at finite temperature in two-color QCD (QC2D), in order to theoretically confirm whether the anomaly enhancement is universal to the number of colors. In what follows, we apply the FRG method to a LSM of QC2D with two flavors ($N_{f}=2$), which was introduced in Ref. Suenaga:2022uqn. This version of the LSM has been developed mainly to investigate the cold and dense properties of QC2D matter extensively Kawaguchi:2023olk; Kawaguchi:2024iaw; Suenaga:2023xwa (for a review, see Suenaga:2025sln).
@@
The FRG, as a mathematical tool, is one of the most promising functional methods in quantum field theory, capable of capturing essential features of strongly interacting quantum systems Berges:2000ew; Dupuis:2020fhh. To the best of our knowledge, there have been no attempts using the FRG method to describe the thermodynamics of QC2D with both mesonic and baryonic degrees of freedom included. In addition to exploring the thermal behavior of the $U(1)_{A}$ anomaly, this paper also aims to establish an FRG-based framework capable of accounting for the thermodynamic properties of two-color QCD in low-energy regimes Fejos:2025nvd.
@@
In QC2D, (anti)diquarks composed of two (anti)quarks are counted as color singlet hadrons. These (anti)diquark baryons can be treated on an equal footing with mesons due to the pseudoreality of the color group $SU(2)$: ${\bm{2}}\simeq\bar{\bm{2}}$. In terms of chiral representation, this property is reflected by an extension of chiral symmetry; in QC2D, $SU(N_{f})_{L}\times SU(N_{f})_{R}$ chiral symmetry is extended to an $SU(2N_{f})$ one, where (anti)diquark baryons and mesons belong to the same multiplet Kogut:1999iv; Kogut:2000ek. This extended symmetry is often referred to as the Pauli-Gürsey $SU(2N_{f})$ symmetry Pauli:1957voo; Gursey:1958fzy.
@@
Since the LSM in QC2D is constructed upon the Pauli-Gürsey symmetry, one can naturally incorporate baryonic contributions in addition to mesonic ones in evaluating the $U(1)_{A}$ anomaly effects. The diquark contributions cannot be easily captured in ordinary three-color QCD, and, thus we will pay particular attention to separating baryonic and mesonic contributions in order to gain insights into fluctuations specific to QC2D.
@@
In-medium properties of QC2D are currently being investigated by first-principles lattice simulations, especially due to the absence of the sign problem at finite baryon chemical potential Boz:2019enj; Buividovich:2020dks; Astrakhantsev:2020tdl; Iida:2024irv; Braguta:2023yhd. Therefore, in the future, one may expect numerical experiments on the lattice, which are capable of testing our current findings. In QC2D, there is no a priori physical point. In the present analysis, we take physical values of the pion mass and the pion decay constant, while $\eta$ mass is considered as a free parameter, in order to investigate the $U(1)_{A}$ anomaly effects in a wider range. We also examine the fate of the anomaly effects as well as hadron mass modifications in a hot medium to provide useful information for future numerical experiments.
@@
This article is organized as follows. In Sec. II, we briefly review the emergence of the Pauli-Gürsey $SU(4)$ symmetry in QC2D with $N_{f}=2$ and introduce our LSM. In Sec. III, our approach, based on the FRG framework, is explained, and RG flow equations for the couplings are derived. Then, in Sec. IV, we present numerical results on the anomaly coefficients, hadron masses, and the topological susceptibility at finite temperature, by solving the flow equations. Section V is devoted to detailed analyses of baryonic contributions. Finally, Sec. VI is devoted to the conclusions.
@@ II Two-flavor LSM in QC2D
In this section, we briefly review the two-flavor LSM in QC2D introduced in Ref. Suenaga:2022uqn.
@@
${\cal L}_{\rm QC_{2}D}$ $=$ $\bar{\psi}i{\vtop{\halign{#\cr\hfil/\hfil\crcr$D$\crcr}}}\psi$ (3)
@@
where in the second line we made use of the pseudorealities $\tau_{c}^{a}=\tau_{c}^{2}(\tau_{c}^{a})^{T}\tau_{c}^{2}$ and $\sigma^{i}=-\sigma^{2}(\sigma^{i})^{T}\sigma^{2}$ Kogut:1999iv; Kogut:2000ek. In Eq. (4), the covariant derivative reads $D_{\mu}={\partial}_{\mu}-ig_{s}{A}_{\mu}^{a}\tau_{c}^{a}$ with the strong coupling constant $g_{s}$ and $\sigma^{\mu}=({\bm{1}},\sigma^{i})$. Intuitively speaking, the second line in Eq. (4) implies that gluons cannot distinguish quarks and antiquarks due to the pseudoreality of the $SU(2)_{c}$ group and they can be described by a single multiplet $\Psi$ collectively. As a result, the Lagrangian with massless quarks possesses the so-called Pauli-Gürsey $SU(4)$ symmetry, as an extension of $SU(2)_{L}\times SU(2)_{R}$ chiral symmetry Pauli:1957voo; Gursey:1958fzy.
Chiral effective models in QC2D must respect the Pauli-Gürsey $SU(4)$ symmetry in order to accurately describe low-energy hadronic excitations. In particular, the LSM is based on the linear representation of this $SU(4)$ symmetry, in which hadron fields are arranged into a $4\times 4$ matrix $\Sigma$. The transformation of $\Sigma$ under the $SU(4)$ group is given by an underlying quark bilinear: $\Sigma_{ij}\sim\Psi^{T}_{j}\sigma^{2}\tau_{c}^{2}\Psi_{i}$ (i,j=1i,j=144). Hence, $\Sigma$ transforms as
@@
where ${\cal S}^{a}$ and ${\cal P}^{a}$ (a=0a=055) are defined through the meson fields
@@
Here, $B$ ($\bar{B}$) and $B^{\prime}$ ($\bar{B}^{\prime}$) are $0^{+}$ and $0^{-}$ (anti)diquark baryons, respectively. We note that mesons ($\bar{q}q$) and baryons ($qq$) are treated in a unified way reflecting the pseudoreality of $SU(2)$ color group: ${\bm{2}}\simeq\bar{\bm{2}}$. In Eq. (6), $4\times 4$ matrices $X^{a}$ are generators belonging to the algebra of $G/H=SU(4)/Sp(4)$:
$X^{a=1,2,3}$ $=$ $\frac{1}{2\sqrt{2}}\left(\begin{array}[]{cc}\tau_{f}^{a}&0\\ 0&(\tau_{f}^{a})^{T}\\ \end{array}\right)\ ,$ (11)
@@
where $\tau_{f}^{a}$ (a=1a=133) are the Pauli matrices acting on flavor space, $D^{a=4}=\tau_{f}^{2}$, and $D_{f}^{a=5}=i\tau_{f}^{2}$. Furthermore, $X^{0}={\bm{1}}_{4\times 4}/(2\sqrt{2})$, and
@@
It should be noted that, in QC2D with $N_{f}=2$, the breaking pattern of chiral symmetry is $SU(4)\to Sp(4)$, leading to appearance of five Nambu-Goldstone bosons: three pions, a diquark baryon, and an antidiquark baryon.
With Eq. (6), the Lagrangian of LSM takes the form of Suenaga:2022uqn; Suenaga:2025sln
@@
where $I_{A}$ is responsible for the $U(1)_{A}$ anomaly, since $\Sigma\overset{U(1)_{A}}{\to}{\rm e}^{-i\theta_{A}}\Sigma{\rm e}^{-i\theta_{A}}={\rm e}^{-2i\theta_{A}}\Sigma$. As for other anomalous terms, we may think of incorporating the KMT determinant $\det\Sigma+\det\Sigma^{\dagger}$ additionally. This, however, is not an independent invariant due to an identity of $\det\Sigma+\det\Sigma^{\dagger}=\frac{1}{4}I_{A}^{2}+\frac{1}{2}I_{2}-\frac{1}{8}I_{1}^{2}$: therefore, it must be omitted from the potential. [^1] In Eq. (22) the terms containing $\lambda_{1}$ and $\lambda_{2}$ are rearranged as $\lambda_{1}I_{1}^{2}+\lambda_{2}I_{2}=\left(\lambda_{1}-\frac{\lambda_{2}}{4}\right)\big({\rm tr}[\Sigma^{\dagger}\Sigma]\big)^{2}+\lambda_{2}{\rm tr}[(\Sigma^{\dagger}\Sigma)^{2}]\ .$ (23) Thus, matching with the standard notation $\bar{\lambda}_{1}\big({\rm tr}[\Sigma^{\dagger}\Sigma]\big)^{2}+\bar{\lambda}_{2}{\rm tr}[(\Sigma^{\dagger}\Sigma)^{2}]$ yields $\bar{\lambda}_{1}=\lambda_{1}-\lambda_{2}/4$ and $\bar{\lambda}_{2}=\lambda_{2}$.
We note that our LSM contains $12$ hadrons in total, listed in Eqs. (7) and (8), for which the Pauli-Gürsey $SU(4)$ symmetry of the potential (22) in the presence of the $U(1)_{A}$ anomaly is properly respected. This structure is mandatory to derive a closed set of flow equations as will be seen in Sec. III. An FRG analysis of QC2D in the absence of the anomaly, including only the $\sigma$, $\pi$, $B$, and $\bar{B}$ as hadrons within the (Polyakov-)quark-meson-diquark model, was done in Strodthoff:2013cua.
@@ III.1 Basics
In this section, we introduce the concept of the FRG method and derive flow equations for the couplings of the LSM to incorporate quantum and thermal fluctuations. We do not intend to give a comprehensive description on the technique; the reader is referred to the following reviews Berges:2000ew; Dupuis:2020fhh.
The core of the FRG formalism lies on the scale-dependent quantum effective action, $\Gamma_{k}$, which, unlike the ordinary quantum effective action $\Gamma$, incorporates fluctuations only with momenta $q\gtrsim k$, where $k$ plays the role of a scale separation variable. This construction is achieved by introducing a momentum-dependent mass term into the classical action, usually called the regulator, which gives a large mass to the IR modes, effectively freezing and, thus, excluding them from the path integral. As $k\rightarrow 0$, one demands that the regulator vanishes, leading to $\Gamma_{k=0}=\Gamma$, while for $k\rightarrow\infty$ (or practically some UV cutoff $\Lambda_{\rm UV}$) requires that the regulator tends to infinity for all modes, realizing that $\Gamma_{k\rightarrow\infty}=S$, $S$ being the classical action, upon which our quantum field theory is built.
The scale-dependent effective action, $\Gamma_{k}$ obeys the Wetterich equation Wetterich:1992yh:
@@
$\tilde{\partial}_{k}=\partial_{k}R_{k}\frac{\partial}{\partial R_{k}},$ (25)
@@
where once again, both the Tr and Log operations need to be taken in the functional and matrix senses. One may also say that $\tilde{\partial}_{k}$ is a differential operator that acts on only the $k$ dependence of $R_{k}$.
The right-hand side (rhs) of Eq. (24) or (26) is generally nonlocal since $\Gamma_{k}^{\prime\prime}$ remains as a function of ${\cal S}^{a}(x)$ and ${\cal P}^{a}(x)$. In the present analysis, however, we employ the local potential approximation (LPA) and ignore nonlocalities. In the LPA, the scale-dependent effective action is approximated as ($\int_{x}\equiv\int d^{4}x_{E}$)
$\Gamma_{k}=\int_{x}\Big[{\rm tr}[\partial_{i}\Sigma^{\dagger}\partial_{i}\Sigma]+\bar{c}{\rm tr}[\zeta^{\dagger}\Sigma+\Sigma^{\dagger}\zeta]+V_{k}\Big],$ (27)
where $V_{k}$ is now a local function, called the scale-dependent effective potential. Note that, we have adopted the Euclidean signature compared to Eq. (18). Also note tha the explicit symmetry-breaking term does not depend on the separation scale $k$, which follows from the fact that couplings corresponding to linear terms in the field variable do not renormalize. By employing the LPA ansatz (27), what we need to establish is a flow equation for $V_{k}$.
@@
$\Gamma_{k}^{\prime\prime}(p,q)=\big(p^{2}+V_{k}^{\prime\prime}\big)\delta(p+q),$ (28)
where the momentum variables contain both the frequencies and the three-momenta: therefore, in accordance, $p^{2}=\omega_{n}^{2}+{\bm{p}}^{2}$ with $\omega_{n}=2\pi nT$ ($n\in{\mathbb{Z}}$) being the Matsubara frequencies. In the present analysis, we adopt the so-called 3D Litim regulator Litim:2001up, which, in Fourier space, depends only on the three-momentum:
@@
where now both the trace and log operations need to be considered only in the matrix sense, and the rhs includes one single-momentum integral only. Note that, owing to the LPA, wave function renormalizations of all field variables are omitted. The summation over $n$ stands for the infinite sum of the Matsubara frequencies $\omega_{n}$, explicitly showing that we are working at some finite temperature $T$. At the UV scale $k=\Lambda_{\rm UV}$, and $V_{k=\Lambda_{\rm UV}}$ is reduced to the classical potential (22), while $V_{k=0}$ becomes the full quantum effective potential. In solving the flow equation, we need to integrate from $k=\Lambda_{\rm UV}$ to $k=0$ using the initial condition dictated by Eq. (22).